OpenStudy (anonymous):

The nearest star to our solar system is 4.29 light year away. How much is this distance in terms of parsecs? how much parallax would this star show then viewd from two locations of the earth six months apart in its orbit around the sun?

OpenStudy (anonymous):

@satellite73 @ajprincess @UnkleRhaukus @physicsme @phi @Callisto @campbell_st @quarkine @siddhantsharan @Shane_B @SmoothMath

OpenStudy (shane_b):

To start off, 1 light year = 0.30659458 Parsecs

OpenStudy (anonymous):

Distance of the star from the solar system = 4.29 ly 1 light year is the distance travelled by light in one year. 1 light year = Speed of light × 1 year = 3 × 108 × 365 × 24 × 60 × 60 = 94608 × 1011 m ∴4.29 ly = 405868.32 × 1011 m 1 parsec = 3.08 × 1016 m ∴4.29 ly = = 1.32 parsec Using the relation, But, 1 sec = 4.85 × 10–6 rad ∴

OpenStudy (shane_b):

I know that the relation is:\[d=\frac{1}{parallax}\]Where distance is in parsecs and parallax is in arcseconds. I'm trying to figure out how the 6 months plays into this...

OpenStudy (shane_b):

Nevermind, I don't think it actually does.

OpenStudy (shane_b):

So I'd say it's \[\frac{1}{1.32 pc}=0.758~arcseconds\]